A ball of mass 100 gram falls from a height of 5 m and rebounds to a height of 1.25m. Calculate the impulse and average force between the balls and the ground, if the time during which they are in contact.

Grade:

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
Impulse =mv2-mv1 vectorially
v^2/2g=height
=>v=sqrt(2gh)
=>v1=sqrt(100)
=10v2
=sqrt(2gh)
=sqrt(2*10*1.25)=25=5
=>Impulse=0.1 *(5i-(-10i))=1.5 i
Since both velocity are in opposite direction.
F(average)=Impulse*time in contact
time in contact is not provided in the question

Thanks & Regards
Arun Kumar
Askiitians Faculty
IIT Delhi